x power4 plus (2minus x)power4 is equal to 34.


XPOWER4 + (2-X)POWER4 = 34

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given x4+(2-x)4=34
Let AM of x,x-2 i.e x-1=y so
(y+1)4+(y-1)4=34
Expanding this and after cancellation we get
y4+6y2-16=0
let y2=u
So u2+6u-16=0
we getu=-8,2
So y=\pm i 2\sqrt{2},\pm\sqrt{2}
So x=1\pm i2\sqrt{2},1\pm\pmi2\sqrt{2}